Order: KCL 60mEq added to each 1000 mL IV fluid        Supply: KCL 2 mEq/mL         ...

  1. Order: KCL 60mEq added to each 1000 mL IV fluid

       Supply: KCL 2 mEq/mL

         Give: _________mL(s) (Round to the nearest whole number)
Answer #1

ANSWER.

* Calculation of ml of drug required.

      Volume(ml) = Desired dose(D)/Dose in hand(H) * Quantity(Q).

Here

   Desired dose(D) = 60 mEq.

   Dose in hand(H) = 2mEq.

      Quantity(Q) = 1 ml.

As per above formula

   Volume(ml) of drug required = D/H * Q

                                                 = 60/2 * 1

                                                 = 30 ml.
Answer #2
D×V/A =60×1000/2 =30000ml
source: Nursing
answered by: Gajendra